OpenStudy (dmezzullo):

Solve for x: the quantity of x plus 4 over 3 = 2 x = −2 x = 2 x = 2 over 3 x = negative 10 over 3

OpenStudy (dmezzullo):

@jhonyy9

OpenStudy (anonymous):

x = 2

OpenStudy (dmezzullo):

thts wat i figured forgot to put it

jhonyy9 (jhonyy9):

(x+4)/3 =2 x+4=6 x=6-4=2 ok ?